I need to find the prime factorization of 7,623. Which I found that 3*3*7*11*11 is the prime factorization, I just need someone to show me the steps to getting 3*3*7*11*11
Thank You

Answer by CharlesG2(834)

I need to find the prime factorization of 7,623. Which I found that 3*3*7*11*11 is the prime factorization, I just need someone to show me the steps to getting 3*3*7*11*11


7623
7 + 6 + 2 + 3 = 13 + 5 = 18, 18 is divisible by 3, so 3 is factor
7623 = 3 * 2541
2 + 5 + 4 + 1 = 7 + 5 = 12, 12 is divisible by 3, so 3 is a factor again
7623 = 3 * 3 * 847
8 + 4 + 7 = 12 + 7 = 19, 19 is not divisible by 3, 3 no longer a factor
847's last digit not a 5 or a 0, so 5 is not a factor
try 7
84 is divisible by 7, so 840 is divisible by 7, so 847 is divisible by 7
7623 = 3 * 3 * 7 * 121
121 is 11^2
7623 = 3 * 3 * 7 * 11 * 11